# Cardinalities in proofreading

### v1

@@ Abstract
In 1975, G. Pólya pointed out that if two proofreaders found $a$, resp. $b$, errors in a text, and both found $c$ errors, then the unknown number $e$ of all errors is approximated by $e\approx\frac{ab}{c}$. We devise a realistic combinatorial model of Pólya’s result and estimate its performance by the Probabilistic Method.
1 Introduction. The $(p,n)$-proofreader
@@
In [8], see also [1, 3, 4, 7], G. Pólya devised a conjuring trick. Suppose that a proofreader finds $a$ errors in a text, that another proofreader finds $b$ errors, and that the number of errors detected by both is $c$. How can one guess, with these data, the unknown number $e$ ($>0$) of all errors? Well, $a\approx pe$ and $b\approx qe$, where $p$ and $q$ ($\in(0,1]$) are probabilities of error detection for the proofreaders. Also, $c\approx pqe$ due to their independence. Then
$\frac{ab}{c}\approx\frac{peqe}{pqe}=e\,.$
Pólya’s trick, brilliant as it is, is not a rigorous mathematical theorem. In our article, we elaborate it in a combinatorial theorem. We state this Theorem 1.3 below and prove it in the next section. The last third section contains comments on probabilistic vs. discrete models. For $n$ in $\mathbb{N}$ (=$=\{1,2,\dots\}$) we set $[n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
[…]
@@
We denote by $\|\overline{b}\|=\sum_{i=1}^{n}b_{i}$ the number of $1$s in $\overline{b}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) its cardinality (number of elements), for example $|B_{n}|=2^{n}$. If $\overline{s}$ is a finite tuple, then $|\overline{s}|$ denotes the number of terms in it. By $\mathbb{Q}$ we denote the set of fractions.
Let $n\in\mathbb{N}$ and $p=\frac{k}{l}\in(0,1]\cap\mathbb{Q}$, with $k,l\in\mathbb{N}$, $k\leq l$ and coprime $k,l$. We always assume that $p>0$. The $(p,n)$-proofreader is the pair $(P,\overline{s})$ in which
$P=P(\overline{a},\,\overline{b})\colon B_{n}\times B_{n}\to B_{n},\ P(\overline{a},\,\overline{b})=(a_{1}b_{1},\,a_{2}b_{2},\,\dots,\,a_{n}b_{n})\,,$
and
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N}),\ N=l^{n}\,\text{ and }\,\overline{s}_{j}\in B_{n}\,,$
is a uniquely determined $N$-tuple defined as follows. We take the map $f\colon[l]\to\{0,1\}$, $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$, and extend it to
$f\colon[l]^{n}\to B_{n},\ f((a_{1},\,a_{2},\,\dots,\,a_{n}))=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
Then we take the lexicographic ordering
$(\overline{u}_{1},\,\overline{u}_{2},\,\dots,\,\overline{u}_{N})$
of $[l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$. Any ordering of $[l]^{n}$ would do, but we want that $(p,n)$-proofreaders be unique objects. Finally, we set
@@
Let $(P,\overline{s})$ be the $(p,n)$-proofreader and $\overline{t}\in B_{n}$ be a text, in which the errors are positions of $1$s. Then for $j\in[N]=[l^{n}]$, where $l$ is the denominator of $p$, the proofreader finds in the text $\overline{t}$ in the $j$-th run $\overline{s}_{j}$ exactly $\|P(\overline{s}_{j},\overline{t})\|$ errors. We show that proofreaders work in the desired way.
@@
Let $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\ \ (\in B_{n}^{N})\,,$
be the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
$N^{-1}\cdot\big{|}\big{\{}j\in[N]\colon\;i\in I\Rightarrow s_{j,\,i}=1\big{\}}\big{|}=p^{|I|}\,.$
Proof. Let $\overline{s}$ and $I$ be as stated. It is easy to see from the definition of $\overline{s}$ that the number of $j\in[N]$ such that $s_{j,i}=1$ for every $i\in I$ is $k^{|I|}\cdot l^{n-|I|}$. But then
$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}l^{n-|I|}}{l^{n}}=\bigg{(}\frac{k}{l}\bigg{)}^{|I|}=p^{|I|}\,.$
@@
Proofreading of a text is usually done only once, or a few times. We think of the $N$ runs $\overline{s}_{j}$ as of performances of the proofreader in possible worlds ([9]).
We need two independent proofreaders. Let $n\in\mathbb{N}$, $p,q\in(0,1]\cap\mathbb{Q}$ and let $(P,\overline{s})$ and $(P,\overline{u})$ be the $(p,n)$-proofreader and the $(q,n)$-proofreader, respectively, with $|\overline{s}|=M$ and $|\overline{u}|=N$. Two (independent) $(p,q,n)$-proofreaders is the pair $(P,\overline{S})$ such that
@@
$=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots$
$\dots,\,(\overline{s}_{2},\,\overline{u}_{N}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2})\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))$
($\in(B_{n}^{2})^{MN}$). Again, any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we want that pairs of proofreaders be unique. For any $j\in[MN]$ we write
@@
We show that two proofreaders work in the desired way.
@@
Let $(P,\overline{S})$, where
be two $(p,q,n)$-proofreaders. The following holds.
@@
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=1\big{\}}\big{|}=p^{|I|}\,.}$
@@
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow u_{w_{j},\,i}=1\big{\}}\big{|}=q^{|I|}\,.}$
@@
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=u_{w_{j},\,i}=1\big{\}}\big{|}=(pq)^{|I|}\,.}$
Proof. Let $\overline{S}$ and $I$ be as stated. 1. Let $r$ be the number of $j\in[M]$ such that $s_{j,i}=1$ for every $i\in I$. By Proposition 1.1, $\frac{r}{M}=p^{|I|}$. Thus the stated product equals
${\textstyle\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.}$
2. The argument is symmetric, with the number $s$ of $j^{\prime}\in[N]$ such that $u_{j^{\prime},i}=1$ for every $i\in I$. 3. Now the stated product equals
${\textstyle\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.}$
@@
We state our main result, a combinatorial version of Pólya’s trick.
@@
${\textstyle c(p,\,q)=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\,.}$
Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have for every number $a\in[0,m]$ that
@@
${\textstyle\geq 1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
For $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction as +$\infty$+ ∞.
Pólya’s estimate has distance at most $a$ from the number of errors for at least the fraction $1-\dots$ of pairs $(\overline{s}_{v_{j}},\overline{u}_{w_{j}})$ of runs of the two proofreaders. Next, in one extreme regime of Theorem 1.3 we minimize the bound $a$ on the distance. In the other extreme we maximize the fraction of pairs of runs for which the distance is still reasonably small. Deduction of both corollaries is immediate.
@@
In the setup of Theorem 1.3, if $\omega(m)\colon\mathbb{N}\to\mathbb{N}$ is any function such that $\omega(m)\leq\sqrt{m}$ and $\omega(m)\to+\infty$, then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
@@
$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to+\infty)\,.$
We prove Theorem 1.3 next.
@@
We deduce Theorem 1.3 from two propositions.
@@
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
An analogous bound holds when $\overline{s}_{v_{j}}$ replaced by $\overline{u}_{w_{j}}$, and $pm$ by $qm$.
Proposition 2.2
Suppose that $(P,\overline{S})$, where
@@
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
We prove both propositions at the conclusion of this section. Now we deduce from them Theorem 1.3.
Deduction of Theorem 1.3 from Propositions 2.1 and 2.2. Suppose that $(P,\overline{S})$, where
are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$. We set
${\textstyle b=c(p,\,q)\cdot a=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.}$
For $j\in[MN]$, let $A$ be the set of $j$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
@@
$|R-m|:=\big{|}\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big{|}\leq a$
($\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=0$). Thus if $j\in A$, resp. $j\in B$, then
[…]
@@
${\textstyle\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.}$
@@
$\leq$ $2b{\textstyle\big{(}\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big{)}=a\,,}$
and hence $D\supset A\cap B\cap C$. By Propositions 2.1 and 2.2,
${\textstyle\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
$\Box$
We develop tools for proofs of Propositions 2.1 and 2.2. Let $(P,\overline{S})$, where
@@
$P=P(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,,$
in which $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $P$-events
$A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;u_{w_{j},i}=1\}$
@@
$C_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=u_{w_{j},i}=1\}\,.$
By Propositions 1.1 and 1.2, each of these three $n$-tuples is independent. Further, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$. Recall that $B^{c}$ denotes the complementary event to the event $B$ in a probability space. For example,
$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=0\}\,.$
Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $P$-events
${\textstyle A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}}$
@@
${\textstyle C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.}$
We use the following key theorem which restates [2, Corollary A.1.7 on p. 325]. It is proven on [2, pp. 321–325].

### v2

@@ Abstract
In 1975, G. Pólya suggested that if two proofreaders found $a$ and $b$ errors in a text, of which $c$ errors were found by both of them, then a reasonable approximation of the unknown number $e$ of all errors is $e\approx ab/c$. We justify this formula by constructing a realistic model of proofreaders and estimating the efficiency of this model with the help of the bound on large deviations in the Probabilistic Method. In conclusion we discuss the distinction between realistic and probabilistic models of problems.
Dedicated to the memory of my mother Blanka Klazarová (1937–2026)
1 Introduction. $(p,n)$-proofreaders
@@
In [7], G. Pólya considered the following problem (see also [1, 3, 4, 6]). Suppose that a proofreader found $a$ errors in a text, that another proofreader found $b$ errors, and that the number of errors detected by both of them is $c$. With these data, how could one approximate the unknown number $e$ of all errors in the text? Pólya’s solution is this. We have $a\approx pe$ and $b\approx qe$, where $p$ and $q$ in $(0,1]$ are the probabilities of error detection for the two proofreaders. Their independence gives $c\approx pqe$. Thus we get the approximation
$e=\frac{peqe}{pqe}\approx\frac{ab}{c}\,.$
Pólya’s interesting argument is not supported by any rigorous theorem. In this article we provide one. It is Theorem 1.5 which we prove in Section 2. In Section 3 we discuss the two kinds of mathematical models one can devise to solve problems: realistic models and probabilistic models.
For $n$ in $\mathbb{N}=\{1,2,\dots\}$ we set $n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
$\overline{b}=(\overline{b}_{1},\,\overline{b}_{2},\,\dots,\,\overline{b}_{n}),\ \overline{b}_{i}\in\{0,\,1\}\,.$
We denote by
$\|\overline{b}\|=\sum_{i=1}^{n}\overline{b}_{i}$}
@@
the number of $1$s in $\overline{b}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) the number of its elements. For example, $|B_{n}|=2^{n}$. For a tuple $\overline{s}$ we denote by $|\overline{s}|$ its length, the number of terms in $\overline{s}$. By $\mathbb{Q}$ we denote the set of fractions.
Let $p=k/l$ be in $(0,1]\cap\mathbb{Q}$, with $\in\mathbb{N}, klk\leq l$ and coprime $k,l$, and let $n\in\mathbb{N}$. We define the key discrete structure of our article.
Definition 1.1 ($(p,n)$-proofreaders)
$(p,n)$-proofreader $\mathcal{P}$ is the pair $\mathcal{P}=(P,\overline{s})$ with the following components which are uniquely determined by $p$ and $n$.
  1. 1.
    $P=P(\overline{a},\overline{b})\colon B_{n}\times B_{n}\to B_{n}$ is a map given by
    $P(\overline{a},\,\overline{b})=(\overline{a}_{1}\overline{b}_{1},\,\overline{a}_{2}\overline{b}_{2},\,\dots,\,\overline{a}_{n}\overline{b}_{n})\,.$
  2. 2.
    The $N$-tuple of $n$-tuples
    $\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,$,
    where $\overline{s}_{j}\in B_{n}$ and $N=l^{n}$, is defined as follows. We begin with the map $f\colon[l]\to\{0,1\}$ given by $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$. We extend it to the map $f\colon[l]^{n}\to B_{n}$ by
    $f(a_{1},\,a_{2},\,\dots,\,a_{n})=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
    We take the lexicographic ordering $\overline{u}_{1}$, $\overline{u}_{2}$, $\dots$, $\overline{u}_{N}$ of $l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$. Finally, we set
@@
    We call the $N$-tuple $\overline{s}$ the runs of $\mathcal{P}$.
Any ordering of $l]^{n}$ would do. We chose the lexicographic one just concreteness sake.
We show that the proofreaders introduced in Definition 1.1 work according to the parameters $p$ and $n$. The parameter $p$ is the efficiency (probability of error detection) of the proofreader, and the parameter $n$ is the length of the text submitted for proofreading. $\mathcal{P}=(P,\overline{s})$ works in $N$ runs $\overline{s}_{j}$, $j\in[N]$. If $(\overline{s}_{j})_{i}=1$ then in the $j$-th run $\mathcal{P}$ pays attention to the $i$-th position in the text, and if $(\overline{s}_{j})_{i}=0$ then $\mathcal{P}$ ignores it. So if a text $\overline{t}\in B_{n}$ (in which $1$s are errors) is submitted for proofreading, then $\mathcal{P}$ in the run $\overline{s}_{j}$ detects
$\|P(\overline{s}_{j},\,\overline{t})\|\ \ (\in\mathbb{N}_{0})$
errors in $\overline{t}$. The next proposition shows that for every set $I\subset[n]$ of positions in the text, $\mathcal{P}$ pays attention to all positions in II in exactly $p^{|I|}\cdot N$ of the $N$ runs.
@@
Let $\overline{s}=(\overline{s}_{1},\overline{s}_{2},\dots,\overline{s}_{N})$ be the runs of the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
$N^{-1}\cdot\big|\big\{j\in[N]\colon\;\text{$(\overline{s}_{j})_{i}=1$ for every $i\in I$}\}\big|=p^{|I|}\,.$
Proof. Let $\overline{s}$, $p$, $n$, and $I$ be as stated. It is easy to see from the definition of $\overline{s}$ that the number of indices $j\in[N]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$ equals
$k^{|I|}\cdot l^{n-|I|}$
because it is the number of words of length $n$ over the alphabet $[l]$ that have at every position $i\in I$ a letter from $[k]$ ($\subset[l]$) . But then
$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}\cdot l^{n-|I|}}{l^{n}}=\bigg(\frac{k}{l}\bigg)^{|I|}=p^{|I|}\,.$
@@
For Pólya’s argument we need two independent proofreaders. Suppose that $p,q\in(0,1]\cap\mathbb{Q}$ and that $n\in\mathbb{N}$.
Definition 1.3 (two $(p,q,n)$-proofreaders)
Two $(p,q,n)$-proofreaders is the pair $\mathcal{R}=(P,\overline{S})$ with the following components which are uniquely determined by $p$, $q$ and $n$.
  1. 1.
    The map $P$ is the same as in Definition 1.1.
  2. 2.
    We construct the $MN$-tuple $\overline{S}$ of pairs of $n$-tuples as follows. Let $(P,\overline{s})$ be the $(p,n)$-proofreader, $(P,\overline{u})$ be the $(q,n)$-proofreader, and let $M=|\overline{s}|$ and $N=|\overline{u}|$. We set
@@
    $=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{2},\,\overline{u}_{N})\,$,
    $\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\ \ (\in(B_{n}\times B_{n})^{MN})\,.$
    We call the $MN$-tuple $\overline{S}$ the runs of $\mathcal{R}$.
Any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we again want to have a uniquely defined object. For any $j\in[MN]$ we write
@@
In the next proposition, we show that for every set $I\subset[n]$ of positions in the text submitted for proofreading, the two proofreaders pay attention to these positions mutually independently.
@@
Let $\overline{S}=((\overline{s}_{1},\overline{u}_{1}),(\overline{s}_{1},\overline{u}_{2}),\dots,\dots,(\overline{s}_{M},\,\overline{u}_{N}))$ be the runs of two $(p,q,n)$-proofreaders. The following holds.
@@
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=p^{|I|}\,.$}
@@
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=q^{|I|}\,.$}
@@
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=(pq)^{|I|}\,.$}
Proof. Let $\overline{S}$ and $I$ be as stated.
1. Let $r$ be the number of indices $j\in[M]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$. By Proposition 1.2, $\frac{r}{M}=p^{|I|}$. Thus the displayed product equals
$\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.$}
2. The argument is symmetric, with the number $s$ of indices $j\in[N]$ such that $(\overline{u}_{j})_{i}=1$ for every $i\in I$.
3. Now the displayed product equals
$\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.$}
@@ 3. Now the displayed product equals
The following theorem is one of our main results.
@@
$c(p,\,q)=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\ \ (\in(0,\,\frac{1}{8}])\,.$}
Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors and for every real number $a\in[0,m]$ we have
@@
${\textstyle\geq 1-6\exp\big(-\frac{2c(p,\,q)^{2}a^{2}}{m}\big)\,,}$
where for $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction $\frac{\cdots}{0}$ as $\infty$.
The theorem says that the distance between Pólya’s approximation and the actual number $m$ of errors is at most $a$ for at least the displayed fraction 11-$\dots$ of the $MN$ runs of $\mathcal{R}$.
The theorem has two corollaries. We minimize the distance of the approximation from the actual number of errors. Then we maximize the fraction of runs for which the distance is still reasonably small. The proofs are immediate, and we omit them.
@@
In the setup of Theorem 1.5, if $\omega\colon\mathbb{N}\to\mathbb{N}$ is a function such that
[…]
then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
@@
$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to\infty)\,.$
We prove Theorem 1.5 in the next section.
@@
We deduce Theorem 1.5 from the next proposition.
@@
are the two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors the following holds.
  1. 1.
    We have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  2. 2.
    We have
    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{u}_{w_{j}},\,\overline{t})\|-qm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  3. 3.
    We have
    ${\textstyle\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big|\leq a\big\}\big|}$
    ${\textstyle\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
We prove it at the end of this section. Now we deduce Theorem 1.5.
Deduction of Theorem 1.5 from Proposition 2.1. Suppose that $(P,\overline{S})$, where
@@
are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$ is a real number. We set
$b=c(p,\,q)\cdot a=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.$}
Let $A$ be the set of $j$ ($\in[MN])$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
@@
$|R-m|:=\big|\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big|\leq a\,$,
where $\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$. Thus if $j\in A$, resp. $j\in B$, then
$\text{$\|P(\overline{s}_{v_{j}},\,\overline{t})\|=(1+\delta_{1})pm$, resp. $\|P(\overline{u}_{w_{j}},\,\overline{t})\|=(1+\delta_{2})qm$}\,$,
@@
$\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.$}
@@
$\leq$ $2b{\textstyle\big(\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big)=a}$
and therefore $D\supset A\cap B\cap C$. By items 1–3 of Proposition 2.1,
$\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})$}
(for $X\subset[MN]$ we write $X^{c}$ for the complement $MN]\setminus X$). $\Box$
We develop tools for the proof of Proposition 2.1. Let $(P,\overline{S})$, where
@@
$\mathrm{PS}(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,$,
where $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $\mathrm{PS}(MN)$-events
$A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;(\overline{u}_{w_{j}})_{i}=1\}$
@@
$C_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1\}\,.$
By items 1–3 of Proposition 1.4, each of these three $n$-tuples consists of independent events. Also, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$. Recall that $B^{c}$ denotes the complementary event to the event $B$. For example,
$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=0\}\,.$
Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $\mathrm{PS}(MN)$-events
$A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}$}
@@
$C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.$}
We employ the following theorem which follows from [2, Corollary A.1.7 on p. 325] which is proven in [2, pp. 321–325].